If A=(1,2) and B=(4,5).the eqyequat of circle having AB as diameter?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
If AB is diameter then centre = (5/2, 7/2)
 
Hence radius = 3 sqrt(3) /2
 
Hence equation of circle =
 
(x-5/2)² + (y - 7/2)² = 27/4
 
Hope it helps